Please help answer my probability questions!

1. A 4-volume numbered set of books is placed randomly on a shelf. What is the probability that the books will be numbered in the correct order from left to right?

2. Pizza House offers 5 different salads, 4 different kinds of pizza, and 5 different desserts. How many different three-course meals can be ordered?

3. If the probability of a child being a girl is 1/2 and a family plans to have 5 children, what are the odds against having all girls?
Answer #1

1) number of ways to arrange them =4! =4*3*2*1=24

therefore probability that the books will be numbered in the correct order from left to right =1/24

2)

number of different three-course meals =5*4*5 =100

3)

P(all girls) =(1/2)^5 =1/32

therefore odds against having all girls =(1-p)/p =(1-1/32)/(1/32) =31 to 1